OpenStudy (anonymous):

what is the cube root of -343 or does it not exist?

OpenStudy (whpalmer4):

\[\sqrt[3]{-343} = \sqrt[3]{-1*343} = \sqrt[3]{-1}*\sqrt[3]{343}\] Do you know the cube root of 343?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its -7